what is 2x to the 4th y to the -4 z to the -3 over 3x to the 2nd y to the -3 z to the 4th ?
Found 2 solutions by stanbon, dabanfield:
Answer by stanbon(75887)

what is 2x to the 4th y to the -4 z to the -3 over 3x to the 2nd y to the -3 z to the 4th ?
------------------------------------
[2x^4*y-4*z*-3)/(3x^2*y^-3*z^4)
---
= (2/3)(x^7)(1/y)(1/z^7)
---
= (2/3)(x^/yz^7)
---------
Cheers,
Stan H.

Answer by dabanfield(803)

what is 2x to the 4th y to the -4 z to the -3 over 3x to the 2nd y to the -3 z to the 4th ?
(2x^4*y^(-4)*z^(-3))/(3x^2*y^(-3)*z^4) =
(2/3)[((x^4)*(x^(-2))*((y^(-4))*(y^(+3))*((z^(-3)*z^(-4))] =
(2/3)[(x^2)*(y^(-1))*(z^(-7))]